# Infinite-dimensional next-generation reservoir computing

### v2

@@ II.4 The NG-RC methodology
More explicitly, assume that we have a collection of $n$ inputs and outputs $\{(\mathbf{z}_{1},y_{1}),\ldots,(\mathbf{z}_{n},y_{n})\mid\mathbf{z}_{i}\in\mathbb{R}^{d},y_{i}\in\mathbb{R}\}$. For each $t\in\{\tau,\ldots,n\}$, construct the $t$-th $\tau$-delay vector
$\underline{\mathbf{z}_{t}}^{\tau}\coloneqq\left(\begin{pmatrix}z_{t-\tau+1,1}\\ \vdots\\ z_{t-\tau+1,d}\end{pmatrix},\ldots,\begin{pmatrix}z_{t,1}\\ \vdots\\ z_{t,d}\end{pmatrix}\right)\in\left(\mathbb{R}^{d}\right)^{\tau}.$
@@
write the vector containing all $k$-degree monomials of elements of $\underline{\mathbf{z}_{t}}^{\tau}$ as
$\left[\underline{\mathbf{z}_{t}}^{\tau}\right]^{k}$ $\coloneqq\left(\prod_{s=t}^{t-\tau+1}\prod_{u=1}^{d}z_{s,u}^{k_{s,u}}\right)_{\sum_{s,u}k_{s,u}=k,\hskip 0.70004ptk_{s,u}\geq 0}$
$=\left(\prod_{s=1}^{\tau d}z_{t,s}^{k_{t,s}}\right)_{\begin{subarray}{c}\sum_{s=1}^{\tau d}k_{t,s}=k,\\ k_{t,1},\ldots k_{t,\tau d}\geq 0\end{subarray}}.$
@@
NG-RC proposes as link between inputs and outputs the solution of the nonlinear regression that uses the components of $\Phi$ as covariates, that is, it requires solving the optimization problem:
$\operatorname*{\arg\min}_{\mathbf{w}\in\mathbb{R}^{N}}\left\{\frac{1}{n-\tau+1}\sum_{i=\tau}^{n}{\left(\mathbf{w}^{\top}\Phi(\underline{\mathbf{z}_{i}}^{\tau})-y_{i}\right)}^{2}+\lambda\|\mathbf{w}\|^{2}_{2}\right\}.$ (15)
@@
$\mathbf{w}^{*}={(\mathsf{X}^{\top}\mathsf{X}+\lambda_{\textup{reg}}\mathbb{I})}^{-1}\mathsf{X}^{\top}\mathsf{Y},$ (16)
where $\mathbb{I}$ denotes the identity matrix. Consequently, the output $y_{t}\in\mathbb{R}$ of an NG-RC system at each time step $t\in\mathbb{Z}$ corresponding to an input $\underline{\mathbf{z}}\in\left(\mathbb{R}^{d}\right)^{\mathbb{Z}}$ is given by
@@ III Kernelization of NG-RC
$c=\left(\sqrt{\binom{p}{k_{o},k_{1},\ldots,k_{\tau d}}}\right)^{\top}_{\begin{subarray}{c}k_{0}+k_{1}+\ldots+k_{\tau d}=p\\ k_{0},k_{1},\ldots,k_{\tau d}\geq 0\end{subarray}}\in\mathbb{R}^{N}.$ (21)
@@
$\mathbb{H}_{\textup{poly}}=\mathbb{H}_{\textup{NG-RC}}.$ (22)
@@
In the setup of the previous proposition, consider the case $d=1$, $\tau=2$, and $p=2$. In that situation, the two kernels in the previous discussion are given by:
@@
and given that the NG-RC map is given by:
@@
$K^{\textup{NG-RC}}(\underline{\mathbf{z}_{i}}^{\tau},\underline{\mathbf{z}_{j}}^{\tau})=\Phi(\underline{\mathbf{z}_{i}}^{\tau})^{\top}\Phi(\underline{\mathbf{z}_{j}}^{\tau})\\ =1+z_{i}z_{j},z_{i-1}z_{j-1}+z_{i}^{2}z_{j}^{2}+z_{i-1}z_{j-1}z_{i}z_{j}+z_{i-1}^{2}z_{j-1}^{2}.$
@@
Another important observation that is visible in these expressions is the difference in computation complexity between NG-RC and its kernelized version introduced in Proposition 3. Recall that NG-RC produces the vector $\mathbf{w}^{\ast}\in\mathbb{R}^{N}$ in (19) while the polynomial kernel regression yields $\bm{\alpha}^{\ast}\in\mathbb{R}^{n-\tau+1}$. NG-RC is a nonlinear regression on the components of the (six-dimensional in this case) feature map $\Phi$, and to carry it out, all those components have to be evaluated at all the data points; on the contrary, the kernelized version only requires the evaluation of $K^{\textup{poly}}$ at the data points, which is computationally simpler. The kernelized version of NG-RC is, hence, computationally more efficient. This difference is even more visible as $\tau$, $d$, and $p$ increase since the dependence (14) of the number of covariates in the NG-RC regression on those parameters makes them grow rapidly, while the behavior of the computational complexity of the polynomial kernel $K^{\textup{poly}}$ is much more favorable. This difference in computational performance between the two approaches will be more rigorously analyzed later in Section V.2.
@@ IV Infinite-dimensional NG-RC and Volterra kernels
Let $\pi_{t}:\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}\longrightarrow\left(\mathbb{R}^{d}\right)$ such that $\pi_{t}(\underline{\mathbf{z}_{i}})=\mathbf{z}_{i-t}$ be the projection operator onto the $(-t)$-th term of a semi-infinite sequence. Given $\tau\in\mathbb{N}$, define the $\tau$-time-delay operator $T_{\tau}:\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}\longrightarrow\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}$ by $\pi_{t}\left(T_{\tau}(\underline{\mathbf{z}})\right)\coloneqq\pi_{t-\tau}\left(\underline{\mathbf{z}}\right)$ for all $t\in{\mathbb{Z}_{-}}$. Given $M>0$, define the space of $M$-bounded inputs by $K_{M}\coloneqq\{\underline{\mathbf{z}}\in(\mathbb{R}^{d})^{\mathbb{Z}_{-}}\mid\|\pi_{t}\left(\underline{\mathbf{z}}\right)\|_{2}^{2}\leq M,\text{ for all }t\in{\mathbb{Z}_{-}}\}$. Choose $\theta>0$ such that $\theta^{2}M^{2}<1$ and choose some $\lambda$ such that $0<\lambda<\sqrt{1-\theta^{2}M^{2}}$. Define the Volterra kernel $K^{\textup{Volt}}:K_{M}\times K_{M}\longrightarrow\mathbb{R}$ by the recursion
@@
where we notice that the term marked with ()(*)( ∗ ) is the sum of all monomials of up to order $k$ on the variables that appear in the inner products $\mathbf{z}_{i}^{\top}\mathbf{z}_{j},\ldots,\mathbf{z}_{i-\tau+1}^{\top}\mathbf{z}_{j-\tau+1}$. This expression yields the polynomial kernel as a polynomial of some finite degree $p$ on the components of the input terms up to some finite lag $\tau$.
@@
Computation of Volterra Gramians. Even though the Volterra kernel is defined in the space of semi-infinite sequences, in applications, only finite samples of size $n$ of the form $\{(\mathbf{z}_{i},y_{i})\}_{i\in\{1,\ldots,n\}}$ are available. In that situation, it is customary to construct semi-infinite inputs of the form $\underline{\mathbf{z}_{i}}=(\ldots,0,0,\mathbf{z}_{1},\ldots,\mathbf{z}_{i-1},\mathbf{z}_{i})\in(\mathbb{R}^{d})^{\mathbb{Z}_{-}}$, for each $i\in\{1,...,n\}$, and we then define for that sample the Volterra Gram matrix $\mathsf{K}^{\textup{Volt}}\in\mathbb{R}^{n\times n}$ as
@@
$\mathsf{K}_{i,j}=1+\frac{\lambda^{2}\mathsf{K}^{\textup{Volt}}_{i-1,j-1}}{1-\theta^{2}\left<\mathbf{z}_{i},\mathbf{z}_{j}\right>},$ (25)
@@
For a newly available set of inputs $\mathbf{z}_{n+1},...,\mathbf{z}_{n+h}$, the estimator can be used to forecast outputs $\widehat{y}_{n+1},\ldots,\widehat{y}_{n+h}$. Extend the Gram matrix to a rectangular matrix $\mathsf{K}^{\textup{Volt}}\in\mathbb{R}^{n,\times n+h}$ by using the recursion
$\mathsf{K}_{i,n+j}=1+\frac{\lambda^{2}\mathsf{K}^{\textup{Volt}}_{i-1,n+j-1}}{1-\theta^{2}\left<\mathbf{z}_{i},\mathbf{z}_{n+j}\right>}$
@@ V.1 Data generating processes and experimental setup
was chosen. A discrete-time dynamical system was derived using Runge-Kutta45 (RK45) numerical integration with time-step 0.005 to simulate a trajectory with 15001 points. The first 5000 points were reserved for training. The remaining points were reserved for testing. By the celebrated Takens’ embedding theorem [45], the Lorenz dynamical system admits a functional that expresses each observation as a function of seven of its past observations. Thus, the Lorenz dynamical system lies within the premise discussed in Section II.2.
@@
The BEKK model is an input/output parametric time series model that is used in financial econometrics in the forecasting of the conditional covariances of returns of stocks traded in the financial markets [12]. We consider $d$ assets and the BEKK(1, 0, 1) model for their log-returns $\mathbf{r}_{t}$ and associated conditional covariances $H_{t}=\textup{Cov}(\mathbf{r}_{t}\mid\mathbf{r}_{t-1},\mathbf{r}_{t-2},\ldots)$ given by
$\mathbf{r}_{t}$ $=H_{t}^{1/2}\mathbf{z}_{t},\quad\mathbf{z}_{t}\sim\textup{IIDN}(0_{d},\mathbb{I}_{d})$
$H_{t}$ $=CC^{\top}+A\mathbf{r}_{t-1}\mathbf{r}_{t-1}^{\top}A^{\top}+BH_{t-1}B^{\top},$
where the input innovations $\mathbf{z}_{t}$ are Gaussian IID, and the output observations are the conditional covariances $H_{t}$. The diagonal BEKK specification is chosen where $C$ is an upper-triangular matrix, and $A$ and $B$ are diagonal matrices of dimension $d$. It is known that there exists a unique stationary and ergodic solution whenever $A_{ii}>0,|B_{ii}|<1$ for $i=1,\ldots,d$, which expresses $H_{t}$ as a highly nonlinear function of its past inputs [6] (as per our premise in Section II.2). Since the covariance matrices are symmetric, we only need to learn the outputs $\mathbf{h}_{t}=\textup{vech}(H_{t})\in\mathbb{R}^{q}$, $q\coloneqq d(d+1)/2$, where the vech operator stacks the columns of a given square matrix from the principal diagonal downwards. An existing dataset from [14] was used with input dimensions of 15 and output dimensions of 120. The dataset consisted of 3760 input and output points, the first 3007 were reserved for training, and the remaining 753 were reserved for testing. Since the output points were very small, to minimize loss of accuracy due to computational truncation errors, the output values were scaled by 1000. The training output data was further normalized so that each dimension would have 0 mean and variance of 1.
@@ V.2 Time complexities
Following [39, page 280], we compute the time complexities for the NG-RC, polynomial kernel regression and Volterra kernel regression. We assume for both kernel regressions, as per the procedure used in the numerical simulations, that the usual Euclidean dot product was used.
@@ V.3 Cross-validation
The range of parameters cross-validated were chosen so that the regularization was performed over the same set of values. As detailed in the previous section, time complexity for NG-RC grows exponentially for larger lag and degree hyperparameters. It was thus impractical to cross-validate over a large range of parameters as with each increase in number of lag or maximum degree of monomials, the computational time would grow exponentially. Thus, only a smaller range of parameters could be cross-validated over. The polynomial kernel was cross-validated over a larger space of the same delay and degree hyperparameters. When cross-validating for the Lorenz system, since Takens’ embedding says only seven lags are needed, a smaller set of lags (up to 10) were cross-validated over. For the rest of the datasets, up to 101 lags were cross-validated over. Such a large range of hyperparameters was possible because by Proposition 3 and Example 4 the polynomial kernel regression uses the same covariates but is faster when more covariates are considered. Finally, mean square error was chosen to be the metric over which the best hyperparameters were chosen.

### v3

@@ II.4 The NG-RC methodology
More explicitly, assume that we have a collection of $n$ inputs and outputs $\{(\mathbf{z}_{1},y_{1}),\ldots,(\mathbf{z}_{n},y_{n})\mid\mathbf{z}_{i}\in\mathbb{R}^{d},y_{i}\in\mathbb{R}\}$. For each $t\in\{\tau,\ldots,n\}$, construct the $t$-th $\tau$-delay vector $\underline{\mathbf{z}_{t}}^{\tau}$, and the vector $\left[\underline{\mathbf{z}_{t}}^{\tau}\right]^{k}$ containing all $k$-degree monomials of elements of $\underline{\mathbf{z}_{t}}^{\tau}$ as follows:
$\underline{\mathbf{z}_{t}}^{\tau}$ $\coloneqq\left(\begin{pmatrix}z_{t-\tau+1,1}\\ \vdots\\ z_{t-\tau+1,d}\end{pmatrix},\ldots,\begin{pmatrix}z_{t,1}\\ \vdots\\ z_{t,d}\end{pmatrix}\right)\in\left(\mathbb{R}^{d}\right)^{\tau},$
$\left[\underline{\mathbf{z}_{t}}^{\tau}\right]^{k}$ $\coloneqq\left(\prod_{s=t}^{t-\tau+1}\prod_{u=1}^{d}z_{s,u}^{k_{s,u}}\right)_{\sum_{s,u}k_{s,u}=k,\hskip 0.70004ptk_{s,u}\geq 0}.$
@@
we can then rewrite $\left[\underline{\mathbf{z}_{t}}^{\tau}\right]^{k}$
$\left[\underline{\mathbf{z}_{t}}^{\tau}\right]^{k}=\left(\prod_{s=1}^{\tau d}z_{t,s}^{k_{t,s}}\right)_{\begin{subarray}{c}\sum_{s=1}^{\tau d}k_{t,s}=k,\\ k_{t,1},\ldots k_{t,\tau d}\geq 0\end{subarray}},$
which we will use in the proof of Proposition 3.
@@
NG-RC ridge regression. NG-RC proposes as link between inputs and outputs the solution of the ridge regression (nonlinear in inputs, linear in covariates) that uses the components of $\Phi$ as covariates, that is, it requires solving the optimization problem:
$\operatorname*{\arg\min}_{\mathbf{w}\in\mathbb{R}^{N}}\left\{\frac{1}{n-\tau+1}\sum_{i=\tau}^{n}{\left(\Phi(\underline{\mathbf{z}_{i}}^{\tau})^{\top}\mathbf{w}-y_{i}\right)}^{2}+\lambda_{\text{reg}}\|\mathbf{w}\|^{2}_{2}\right\}.$ (16)
@@
$\mathbf{w}^{*}={(\mathsf{X}^{\top}\mathsf{X}+\lambda_{\textup{reg}}\mathbb{I}_{N})}^{-1}\mathsf{X}^{\top}\mathsf{Y},$ (17)
where $\mathbb{I}_{N}$ denotes the identity matrix of dimension $N$. Consequently, the output $y_{t}\in\mathbb{R}$ of an NG-RC system at each time step $t\in\mathbb{Z}$ corresponding to an input $\underline{\mathbf{z}}\in\left(\mathbb{R}^{d}\right)^{\mathbb{Z}}$ is given by
@@ III Kernelization of NG-RC
$c=\left(\sqrt{\binom{p}{k_{o},k_{1},\ldots,k_{\tau d}}}\right)^{\top}_{\begin{subarray}{c}k_{0}+k_{1}+\ldots+k_{\tau d}=p\\ k_{0},k_{1},\ldots,k_{\tau d}\geq 0\end{subarray}}\in\mathbb{R}^{N}.$
@@
$\mathbb{H}_{\textup{poly}}=\mathbb{H}_{\textup{NG-RC}}.$
@@
In the setup of the previous proposition, consider the case $d=1$, $\tau=2$, and $p=2$. In that situation, the two kernels in the previous discussion are defined as:
@@
and provided that the NG-RC map is given by:
@@
$K^{\textup{NG-RC}}(\underline{\mathbf{z}_{i}}^{\tau},\underline{\mathbf{z}_{j}}^{\tau})=\Phi(\underline{\mathbf{z}_{i}}^{\tau})^{\top}\Phi(\underline{\mathbf{z}_{j}}^{\tau})\\ =1+z_{i}z_{j}+z_{i-1}z_{j-1}+z_{i}^{2}z_{j}^{2}+z_{i-1}z_{j-1}z_{i}z_{j}+z_{i-1}^{2}z_{j-1}^{2}.$
@@
Another important observation that is visible in these expressions is the difference in computation complexity between NG-RC and its kernelized version introduced in Proposition 3. Recall that NG-RC produces the vector $\mathbf{w}^{\ast}\in\mathbb{R}^{N}$ in (20) while the polynomial kernel regression yields $\bm{\alpha}^{\ast}\in\mathbb{R}^{n-\tau+1}$. NG-RC is a nonlinear (on inputs) regression on the components of the (six-dimensional in this case) feature map $\Phi$, and to carry it out, all those components have to be evaluated at all the data points; on the contrary, the kernelized version only requires the evaluation of $K^{\textup{poly}}$ at the data points, which is computationally simpler. The kernelized version of NG-RC is, hence, computationally more efficient. This difference is even more visible as $\tau$, $d$, and $p$ increase since the dependence (15) of the number of covariates in the NG-RC regression on those parameters makes them grow rapidly, while the behavior of the computational complexity of the polynomial kernel $K^{\textup{poly}}$ is much more favorable. This difference in computational performance between the two approaches will be more rigorously analyzed later in Section V.2.
@@ IV Infinite-dimensional NG-RC and Volterra kernels
Let $\pi_{t}:\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}\longrightarrow\left(\mathbb{R}^{d}\right)$, $t\in\mathbb{Z}_{+}$, such that $\pi_{t}(\underline{\mathbf{z}_{i}})=\mathbf{z}_{i-t}$ be the projection operator onto the $(-t)$-th term of a semi-infinite sequence. Given $\tau\in\mathbb{N}$, define the $\tau$-time-delay operator $T_{\tau}:\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}\longrightarrow\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}$ by $\pi_{t}\left(T_{\tau}(\underline{\mathbf{z}})\right)\coloneqq\pi_{t-\tau}\left(\underline{\mathbf{z}}\right)$ for all $t\in{\mathbb{Z}_{-}}$. Given $M>0$, define the space of $M$-bounded inputs by $K_{M}\coloneqq\{\underline{\mathbf{z}}\in(\mathbb{R}^{d})^{\mathbb{Z}_{-}}\mid\|\pi_{t}\left(\underline{\mathbf{z}}\right)\|_{2}^{2}\leq M,\text{ for all }t\in{\mathbb{Z}_{-}}\}$. Choose $\theta>0$ such that $\theta^{2}M^{2}<1$ and choose some $\lambda$ such that $0<\lambda<\sqrt{1-\theta^{2}M^{2}}$. Define the Volterra kernel $K^{\textup{Volt}}:K_{M}\times K_{M}\longrightarrow\mathbb{R}$ by the recursion
@@
where we notice that the term marked with ()(*)( ∗ ) is the sum of all monomials of order $k$ on the variables that appear in the inner products $\mathbf{z}_{i}^{\top}\mathbf{z}_{j},\ldots,\mathbf{z}_{i-\tau+1}^{\top}\mathbf{z}_{j-\tau+1}$. This expression yields the polynomial kernel as a polynomial of some finite degree $p$ on the components of the input terms up to some finite lag $\tau$.
@@
Computation of Volterra Gramians. Even though the Volterra kernel is defined in the space of semi-infinite sequences, in applications, only finite samples of size $n$ of the form $\{(\mathbf{z}_{i},y_{i})\}_{i\in\{1,\ldots,n\}}$ are available. In that situation, it is customary to construct semi-infinite inputs of the form $\underline{\mathbf{z}_{i}}=(\ldots,0,0,\mathbf{z}_{1},\ldots,\mathbf{z}_{i-1},\mathbf{z}_{i})\in(\mathbb{R}^{d})^{\mathbb{Z}_{-}}$, for each $i\in\{1,\ldots,n\}$, and we then define for that sample the Volterra Gram matrix $\mathsf{K}^{\textup{Volt}}\in\mathbb{R}^{n\times n}$ as
@@
$\mathsf{K}^{\textup{Volt}}_{i,j}=1+\frac{\lambda^{2}\mathsf{K}^{\textup{Volt}}_{i-1,j-1}}{1-\theta^{2}\left<\mathbf{z}_{i},\mathbf{z}_{j}\right>},$ (24)
@@
For a newly available set of inputs $\mathbf{z}_{n+1},\ldots,\mathbf{z}_{n+h}$, the estimator can be used to forecast outputs $\widehat{y}_{n+1},\ldots,\widehat{y}_{n+h}$. Extend the Gram matrix to a rectangular matrix $\mathsf{K}^{\textup{Volt}}\in\mathbb{R}^{n,\times n+h}$ by using the recursion
$\mathsf{K}^{\textup{Volt}}_{i,n+j}=1+\frac{\lambda^{2}\mathsf{K}^{\textup{Volt}}_{i-1,n+j-1}}{1-\theta^{2}\left<\mathbf{z}_{i},\mathbf{z}_{n+j}\right>}$
@@ V.1 Data generating processes and experimental setup
was chosen. A discrete-time dynamical system was derived using Runge-Kutta45 (RK45) numerical integration with time-step 0.005 to simulate a trajectory with 15001 points. The first 5000 points were reserved for training. The remaining points were reserved for testing. Although we consider all coordinates, one could potentially consider reconstructing the Lorenz dynamical system out of partial observations. In this case, the $\tau$-delay would be especially important due to the celebrated Takens’ embedding theorem [36]. Regardless, even in the fully observable case used in this paper, the Lorenz dynamical system lies within the premise discussed in Section II.2.
@@
The BEKK model is an input/output parametric time series model that is used in financial econometrics in the forecasting of the conditional covariances of returns of stocks traded in the financial markets [47]. We consider $d$ assets and the BEKK(1, 0, 1) model for their log-returns $\mathbf{r}_{t}$ and associated conditional covariances $\Sigma_{t}=\textup{Cov}(\mathbf{r}_{t}\mid\mathbf{r}_{t-1},\mathbf{r}_{t-2},\ldots)$ given by
$\mathbf{r}_{t}$ $=\Sigma_{t}^{1/2}\mathbf{z}_{t},\quad\mathbf{z}_{t}\sim\textup{IIDN}(0_{d},\mathbb{I}_{d})$
$\Sigma_{t}$ $=CC^{\top}+A\mathbf{r}_{t-1}\mathbf{r}_{t-1}^{\top}A^{\top}+B\Sigma_{t-1}B^{\top},$
where the input innovations $\mathbf{z}_{t}$ are Gaussian IID, and the output observations are the conditional covariances $\Sigma_{t}$. The diagonal BEKK specification is chosen where $C$ is an upper-triangular matrix, and $A$ and $B$ are diagonal matrices of dimension $d$. It is known that there exists a unique stationary and ergodic solution whenever $A_{ii}>0,|B_{ii}|<1$ for $i=1,\ldots,d$, which expresses $\Sigma_{t}$ as a highly nonlinear function of its past inputs [48] (as per our premise in Section II.2). Since the covariance matrices are symmetric, we only need to learn the outputs $\mathbf{h}_{t}=\textup{vech}\Sigma_{t})\in\mathbb{R}^{q}$, $q\coloneqq d(d+1)/2$, where the vech operator stacks the columns of a given square matrix from the principal diagonal downwards. An existing dataset from [31] was used with input dimensions of 15 and output dimensions of 120. The dataset consisted of 3760 input and output points, the first 3007 were reserved for training, and the remaining 753 were reserved for testing. Since the output points were very small, to minimize loss of accuracy due to computational truncation errors, the output values were scaled by 1000. The training output data was further normalized so that each dimension would have 0 mean and variance of 1.
@@ V.2 Time complexities
Following [29, page 280], we compute the time complexities for the NG-RC, polynomial kernel regression and Volterra kernel regression. We assume for both kernel regressions, as per the procedure used in the numerical simulations, that the usual Euclidean dot product was used.
@@ V.3 Cross-validation
The range of parameters cross-validated were chosen so that the regularization was performed over the same set of values. As detailed in the previous section, time complexity for NG-RC grows exponentially for larger lag and degree hyperparameters. It was thus impractical to cross-validate over a large range of parameters as with each increase in number of lag or maximum degree of monomials, the computational time would grow exponentially. Thus, only a smaller range of parameters could be cross-validated over. The polynomial kernel was cross-validated over a larger space of the same delay and degree hyperparameters. When cross-validating for the Lorenz system, with fully observable coordinates, only one delay suffices to reconstruct the dynamical system. However, in practice, a higher number of lags may offer superior predictive performance. Hence we allowed for up to 10 lags to cross-validate over in the case of Lorenz. For the rest of the datasets, up to 101 lags were cross-validated over. Such a large range of hyperparameters was possible because by Proposition 3 and Example 4 the polynomial kernel regression uses the same covariates but is faster when more covariates are considered. Finally, mean squared error was chosen to be the metric over which the best hyperparameters were chosen.